OpenStudy (anonymous):

solve for x,y, and z 4x+2y+5z=8 5x+3y+4z=9 6x+3y+3z=3

OpenStudy (kira_yamato):

4x+2y+5z=8 -----(1) 5x+3y+4z=9 -----(2) 6x+3y+3z=3 -----(3) Manipulate equation (3) to get the simplest form and plug that into (1) and (2) first.

OpenStudy (anonymous):

oh

OpenStudy (kira_yamato):

Hope that hint was helpful ^_^

OpenStudy (anonymous):

yeah thanks

OpenStudy (kira_yamato):

No problem. I'll be here if you need me ^_^ By the way, I just solved it with a calculator ^_^